BROADENING AND SHIFT COEFFICIENTS IN THE ^{12}C^{16}O_{2} AND ^{13}C^{16}O_{2} LASER BANDS

\underline{D. CHRIS BENNER}, V. MALATHY DEVI, Department of Physics, The College of William and Mary, Box 8795, Williamsburg, VA 23187-8795; M. A. H. SMITH, C. P. RINSLAND, Atmospheric Sciences, NASA Langley Research Center, Mail Stop 401A, Hampton, VA 23681-2199; L. R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109; and M. DULICK, National Solar Observatory, P. O. Box 26732, Tucson, AZ 85726-6732.

In a previous study^{a} we had reported air- and N_{2}-broadening and pressure-induced shift coefficients for 46 rovibrational transitions in the ^{12}C^{16}O_{2} 00^{0}1-(10^{0}0-02^{0}1)_{I} and 00^{0}1-(10^{0}0-02^{0}1)_{II} laser bands near 10 μm These parameters were determined from spectra recorded with the McMath-Pierce Fourier transform spectrometer (FTS) of the National Solar Observatory on Kitt Peak, Arizona. We now report similar measurements of air- and N_{2}-broadening and shifts for the two 13C^{16}O_{2} laser bands, plus new determinations of self-broadening and shifts in the ^{12}C^{16}O_{2} laser bands. These results were obtained from simultaneous analysis of 30 long-path absorption spectra using a multispectrum nonlinear least-squares technique^{b}. In addition to the 10 spectra of air- and N_{2}-broadened natural CO_{2} analyzed previously, we have included 20 new spectra recorded at room temperature and 0.005 cm^{−1} resolution using the McMath-Pierce FTS. These additional spectra include two low pressure, four air-broadened and four N_{2}-broadened spectra obtained with a 90% ^{13}C-enriched CO_{2} sample and 10 self-broadened spectra obtained with a high purity natural CO_{2} sample. The self-broadened spectra were recorded with pathlengths of 73, 97, 193 and 433 m and CO_{2} pressures varying between 3 Torr and 400 Torr. The 13CO_{2} spectra were obtained with a pathlength of 84 m with sample pressures ¡ 4 Torr for the low-pressure spectra and from 100 to 500 Torr for the air- and N_{2}-broadened mixtures. By combining the spectra of ^{12}CO_{2} and ^{13}CO_{2} in the same fit we were able to obtain a consistent set of line parameters for both molecules. The results obtained for the ^{12}CO_{2} and 13CO_{2} laser bands will be compared with each other, with the values in the HITRAN database^{c}, and with available values reported in the literature.

aV. Malathy Devi, D. Chris Benner, M. A. H. Smith and C. P. Rinsland, JQSRT \underline{59}, 137-149 (1998).

bD. Chris Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith and D. Atkins, JQSRT \underline{53}, 705-721 (1995).

cL. S. Rothman et al., JQSRT \underline{60}, 665-710 (1998).